Can someone explain how to do these three:
0.013x=0.546
x/1200= -0.12
-16.7x=367.4

Answer by haileytucki(390)

First one:
0.01x=0.55
Divide each term in the equation by 0.01.
(0.01x)/(0.01)=(0.55)/(0.01)
Cancel the common factor of 0.01.
x=(0.55)/(0.01)
Divide 0.55 by 0.01 to get 42.
x=42

Second one:
(x)/(1200)=-0.12
Multiply each term in the equation by 1200.
(x)/(1200)*1200=-0.12*1200
Simplify the left-hand side of the equation by canceling the common factors.
x=-0.12*1200
Multiply -0.12 by 1200 to get -144.
x=-144
Third one:
-16.7x=367.4
Divide each term in the equation by -16.7.
-(16.7x)/(-16.7)=(367.4)/(-16.7)
Cancel the common factor of -16.7.
x=(367.4)/(-16.7)
Simplify the right-hand side of the equation by simplifying each term.
x=-22